It has been found from experience that the mean breaking strength of a brand of thread is 500 grams. With a S.D of 40gms. From the supplies, received during the last month, a sample of 16 pieces of thread was tested which showed a mean strength of 450 grams. Can we conclude that the thread supplied is inferior.


1
Expert's answer
2022-02-01T16:18:12-0500

"n=16\\\bar x=450\\\sigma=40"

The null and alternative hypothesis are,

"H_0:\mu=500\\vs\\H_1:\mu\lt500"

The test statistic is given as,

"Z={\bar x-\mu\over{\sigma\sqrt{n}}}={450-500\over{40\over\sqrt{16}}}={-50\over10}=-5"

The critical value is,

"Z_{\alpha}=Z_{0.05}=-1.645"

The null hypothesis is rejected if "Z\lt Z_{0.05}".

Clearly, "Z=-5\lt Z_{0.05}=-1.645" therefore, we reject the null hypothesis and conclude that there is sufficient evidence to indicate that the  thread supplied is inferior at "\alpha=0.05" level of significance.